OpenStudy (anonymous):

If f(x) = x^2 - 1, find f(3). 5 2 8 -2

OpenStudy (phi):

f(3) is short hand for : everywhere you see x, replace it with a 3 in x^2 -1 if you change the x into a 3, what do you get ?

OpenStudy (anonymous):

wait why replace x with 3?

OpenStudy (anonymous):

f(3)= 3x^2

OpenStudy (phi):

f(x) = x^2 - 1 change the x into a 3. If you do that , you can't get f(3)= 3x^2 because the x is not there.

OpenStudy (phi):

wait why replace x with 3? Because that is what f(3) means. you start with f(x) = x^2 - 1 f(3) is "code" for replace (change) the x to a 3 if you have a pencil and paper you would write down x^2 -1, erase the x, and put a 3 in its place

OpenStudy (anonymous):

f(3) = 3^2-1

OpenStudy (phi):

yes, now do the arithmetic. remember that 3^2 is short for 3*3

OpenStudy (anonymous):

f(3)= 9-1

OpenStudy (phi):

yes, and 9-1 is ?

OpenStudy (anonymous):

8 thanks i get it